OpenStudy (anonymous):

Use the binomial theorem to expand (d-4b)^3

OpenStudy (anonymous):

do you know the coefficients when you expand (x + y)^3 ????

OpenStudy (anonymous):

I dont want a staright answer. But I need HELP on finding the answer

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

it's 1-3-3-1... \(\large (x+y)^3 = 1\cdot x^3y^0 + 3\cdot x^2y^1 + 3\cdot x^1y^2 + 1\cdot x^0y^3 \) so in your expression \(\large(d-4b)^3 \), let x=d and y=-4b. just plug it into the formula.

OpenStudy (anonymous):

And the answer is d^3-12bd^2+48b^2d-64b^3. Right?